a) Obtain the equation of the least squares line.
b) Calculate the coefficient of determination.
c) Calculate an estimate of the error standard deviation (sigma) in the simple linear regression model.
Your T For the past decade, rubber powder has been used In asphalt cement to improve performance. An article Includes a (MPa) on x - cube strength (MPa) based on the following sample data regression of y-axial strength x112.3 97.0 92.7 86.0 102.0 99.2 95.8 103.5 89.0 86.7 y| 74.5 71.4 57.6 49.2 74.5 73.4 68.1 59.2 57.7 48.0 (a) Obtain the equation of the least squares line. (Round all numerical values to four decimal places.) Interpret the slope O A one MPa decrease in axial strength is assoclated with an increase in the predicted cube strength equal to the slope. In ax O A one MPa increase in cube strength is associated with an increase in the predicted axial strength equal to the slope. O A one MPa decrease in cube strength is associated with an increase in the predicted axial strength equal to the slope. (b) Calculate the coefficient of determination. (Round your answer to four decimal places)
Answer #1

Following table shows the calculations :

X Y X^2 Y^2 XY
112.3 74.5 12611.29 5550.25 8366.35
97 71.4 9409 5097.96 6925.8
92.7 57.6 8593.29 3317.76 5339.52
86 49.2 7396 2420.64 4231.2
102 74.5 10404 5550.25 7599
99.2 73.4 9840.64 5387.56 7281.28
95.8 68.1 9177.64 4637.61 6523.98
103.5 59.2 10712.25 3504.64 6127.2
89 57.7 7921 3329.29 5135.3
86.7 48 7516.89 2304 4161.6
Total 964.2 633.6 93582 41099.96 61691.23

Sample size: n =  10

Now,

S_{yy}=\sum y^{2}-\frac{\left (\sum y \right )^{2}}{n}=955.064

S_{xx}=\sum x^{2}-\frac{\left (\sum x \right )^{2}}{n}=613.836

S_{xy}=\sum xy-\frac{\left (\sum x \right )\left (\sum y \right )}{n}=599.518

(A)

Slope of the regression equation is

b_{1}=\frac{S_{xy}}{S_{xx}}=0.9767

and intercept of the equation will be

b_{0}=\frac{1}{n}(\sum y - b_{1} \sum x)=-30.8110

So the regression equation will be

y'= -30.8110+0.9767x

(b)

SSR=\frac{S^{2}_{xy}}{S_{xx}}=585.5339738
SSE=S_{yy}-\frac{S^{2}_{xy}}{S_{xx}}=369.5300262

SST=955.064

The coefficient of determination is

R^{2}=\frac{SSR}{SST}=0.6131

The estimate of the error standard deviation (sigma) in the simple linear regression model is

\sigma=\sqrt{\frac{SSE}{n-2}}=\sqrt{\frac{369.5300262}{8}}=6.7964